# BalanceKV: KV Cache Compression through Discrepancy Theory

### v1

@@ 3.1 SoftmaxBalance
$\leq O\left(\sqrt{s}\cdot\log(ns/\delta)\cdot\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\exp\left(\max_{j\in[n]}\frac{\left\|k_{j}\right\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\max_{j\in[n]}\|v_{j}\|_{2}\right).$
The proof of the above theorem uses the breakthrough result of [ALS21] for the vector balancing problem, one of the main problems in discrepancy theory. Given a set of vectors $k_{1},\ldots,k_{n}$ the result of [ALS21] produces a subset $C$ of these vectors of at most half the size such that for any vector $q$ we have that $\sum_{k\in C}\langle k,q\rangle\approx\sum_{k\in[n]\setminus C}\langle k,q\rangle$ with high probability. Our main contribution is an algorithm for the vector balancing problem with respect to the function $\exp(\langle k,\cdot\rangle/\sqrt{d})v$ as compared to $\langle k,\cdot\rangle$ in the case of [ALS21]. A full proof of Theorem 3.2 can be found in Section 3.4.
3.2 MergeAndReduce
Figure 1: Illustration of the tree structure of MergeAndReduce. It depicts the state of the algorithm after the last token in $C^{0}$ has arrived, before updating $C^{0},\ldots,C^{T}$ with SoftmaxBalance. Whenever line 13 in Algorithm 3 is triggered for some $C^{i}$, we mark the next batch on its level of recursion as $C^{i}$, so that before each new token arrives $C^{0},\ldots,C^{T-1}$ form a root to leaf path in one of the binary trees.
As briefly mentioned above in Section 3, MergeAndReduce is a streaming version of SoftmaxBalance. The idea is to partition the stream of tokens into batches of size $t$, split the batches into pairs, apply SoftmaxBalance to the batches, merge the outputs of SoftmaxBalance on the batches in one pair, and then repeat recursively on the new set of batches $T-1$ times if the desired compression rate is $2^{T}$– see Figure 1. We run all the levels of recursion in parallel, which means that at any time step we keep only one (possibly empty) batch per each level of recursion, and we refer to these batches as $C^{0},\ldots,C^{T-1}$. We store the points which survived all the levels of recursion in the memory cell called $C^{T}$.
If we set batch size $t$ to be about $1/\varepsilon$ (see Theorem 3.3 below for the more precise setting), we obtain a streaming algorithm that approximates $\sum_{i=1}^{j}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right)v_{i}$ at any point $j$ in the stream using total space $O^{*}\left(e^{2r^{2}/\sqrt{d}}/\varepsilon\right)$ and runtime $O^{*}\left(e^{4r^{2}/\sqrt{d}}/\varepsilon^{2}\right)$ per step, where $r$ is an upper bound on the norms of key and query embeddings.
As before, an important aspect is that MergeAndReduce can handle value embeddings of dimension not necessarily equal to that of key and query embeddings. Thus, when run on scalars $v_{i}=1$ for all $i$, it can also be used to approximate softmax normalization at any point $j$ in the stream. This is the main subroutine used in BalanceKV to approximate Attn$(q_{j},K_{j},V_{j})$.
@@ 3.2 MergeAndReduce
The formal guarantees are given by
Theorem 3.3.
  • batch size $t=O^{*}(e^{2r^{2}/\sqrt{d}}/{\varepsilon})$
  • compression rate $2^{-T}$ with $T=\log(n/t)$.
Then MergeAndReduce on input parameters $t,r,d,s,\varepsilon$, outputs at every step $j$ of the stream subsets of key-value embedding pairs $C^{0},\ldots,C^{T}\subset C:=\{(k_{1},v_{1}),\ldots,(k_{n},v_{n})\}$ such that,
$z_{j}:=\sum_{i=0}^{T}2^{i}\sum_{\{k,v\}\in C^{i}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v,$
satisfies with probability at least $1-1/\mathrm{poly}(n)$,
The total memory of the algorithm is $O^{*}(e^{2r^{2}/\sqrt{d}}/{\varepsilon})$ and its runtime during the $j$-th iteration is $O^{*}(e^{4r^{2}/\sqrt{d}}/{\varepsilon^{2}})$
@@
Note that since in the statement of the theorem $T=\log_{2}(n/t)$, condition until in line 11 of MergeAndReduce is redundant.
Let us first consider the performance of the procedure at time steps which are multiples of $t$. Observe that at any such $j$-th step the procedure is an online implementation of the following simple offline recursive algorithm on dataset $\{(k_{1},v_{1}),\ldots,(k_{j},v_{j})\}$:
@@
    Set $p=j/t$ and $i=1$. Split the dataset $\{(k_{1},v_{1}),\ldots,(k_{j},v_{j})\}$ into batches $B^{0}_{1},\ldots,B^{0}_{p}$ of size $t$.
@@
    • Update $p\leftarrow\lfloor p/2\rfloor$, $i\leftarrow i+1$.
@@
Observe that after each iteration of step 2 the number of batches decreases by a factor of two. The maximum batch size is always bounded by $t$. This is because a batch $B_{l}$ at iteration $i$ of step 2 is a union of SoftmaxBalance$(B_{2l-1})$ and SoftmaxBalance$(B_{2l})$ for $B_{2l-1}$ and $B_{2l}$ at iteration $i-1$ of step 2, and SoftmaxBalance reduces the size of the dataset which it has been applied to at least by a factor of 2.
The memory of the procedure is the collection of memory cells $C^{i}$, and $|C^{i}|\leq t$. Since at time step $j$ at most $\log_{2}(p)\leq\log_{2}(n/t)$ memory cells are occupied, the total memory is bounded by $O(dt\log_{2}(n/t))=O(dtT),$ which, using the $O^{*}$ notation, is equal to $O^{*}(e^{2r^{2}/\sqrt{d}}/{\varepsilon})$.
@@
$\leq\sum_{i=1}^{T}\left\|2^{i}\sum_{\{k,v\}\in B^{i}\cup C^{i}}\right.\left.\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-2^{i-1}\sum_{\{k,v\}\in B^{i-1}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v\right\|_{2}$
$\leq\sum_{i=1}^{T}2^{i-1}\left\|\sum_{\{k,v\}\in B^{i}\cup C^{i}}\right.\left.\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\in B^{i-1}\backslash(B^{i}\cup C^{i})}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v\right\|_{2}.$
We will refer to the $i$-th summand on the right hand side as the error produced by the $i$-th iteration of step 2. At the $i$-th iteration of step 2 we apply SoftmaxBalance to $p/2^{i-1}$ batches $B^{i-1}_{1},B^{i-1}_{2},\ldots$ of size $t$, and we save the outputs of SoftmaxBalance in batches $C^{i},B^{i}_{1},B^{i}_{2},\ldots$. Therefore, by Theorem 3.2 and triangle inequality, the error vector produced by the procedure at the $i$-th iteration of step 2 has $l_{2}$ norm bounded by
$O\left(2^{i-1}\cdot\sqrt{s}\cdot\log(sn)\cdot\left(\frac{p}{2^{i-1}}\right)\cdot\ e^{r^{2}/\sqrt{d}}\max_{l\in[n]}\|v_{l}\|_{2}\right)=O\left(\sqrt{s}\cdot\log(sn)\cdot\ p\cdot e^{r^{2}/\sqrt{d}}\max_{l\in[n]}\|v_{l}\|_{2}\right),$
@@
$O\left(\sqrt{s}\cdot\log(sn)\cdot T\cdot p\cdot e^{r^{2}/\sqrt{d}}\max_{l\in[n]}\|v_{l}\|_{2}\right).$
By definition, $p=j/t$. In order to ensure that the statement of the theorem is correct, the upper bound on the $l_{2}$ norm of the error vector of the procedure should be less than the desired error $e^{-r^{2}/\sqrt{d}}\cdot\max_{l\in[n]}\|v_{l}\|_{2}$:
$O\left(\sqrt{s}\cdot\log(sn)\cdot T\cdot\frac{j}{t}\cdot e^{r^{2}/\sqrt{d}}\max_{l\in[n]}\|v_{l}\|_{2}\right)\leq\varepsilon j\cdot e^{-r^{2}/\sqrt{d}}\cdot\max_{l\in[n]}\|v_{l}\|_{2}.$
And, if we set
@@
the above inequality holds. In the $O^{*}$ notation, since $s\leq d$, $t=O^{*}(e^{2r^{2}/\sqrt{d}}/{\varepsilon})$.
Runtime during one time step.
At worst, during $j$-th time step the algorithm has to launch SoftmaxBalance $\log_{2}(p)\leq\log_{2}(n/t)=T$ times on batches of size $t$, so the runtime is bounded by $O(dt^{2}T)$. In the $O^{*}$ notation, the runtime is equal to $O^{*}(e^{4r^{2}/\sqrt{d}}/\varepsilon^{2}).$
@@
3.3 Proof of Theorem 3.1 (Theoretical Guarantee of BalanceKV)
Finally equipped with Theorem 3.2 and Theorem 3.3 we now state the proof of Theorem 3.1.
@@
[…]
Performance of the algorithm.
At time step $j$, procedure MR-Denominator returns subsets $K_{1},\ldots,K_{T}$ such that
@@
$\left|\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right)-\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in K^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)\right|\leq\varepsilon\cdot j\cdot e^{-r^{2}/\sqrt{d}}\leq\varepsilon\cdot\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right).$
[…]
therefore we get
3.4 Proof of Theorem 3.2 (Theoretical Guarantee of SoftmaxBalance)
Theorem 3.4 (Theorem 1.1 in [ALS21]).
[…]
2:$c\leftarrow 30\log(n/\delta)$
3:$U_{-},U_{+}\leftarrow\emptyset$
4:for $i$ from 1 and until the end of the stream do
6:      Fail
7:   end if
9:   $\varepsilon_{i}\leftarrow\left\{{\begin{array}[]{ll}+1&\text{with probability}\,p_{i}\\ -1&\text{otherwise}\end{array}}\right.$
10:   $U_{\varepsilon_{i}}\leftarrow U_{\varepsilon_{i}}\cup\{k_{i}\}$
11:end for
12:if $|U_{+}|\leq|U_{-}|$ then
13:   output: $U_{+}$
14:else
15:   output: $U_{-}$
16:end if
Proof of Theorem 3.2.
Define for any $k\in\mathbb{R}^{d}$ an embedding function $\varphi(k):$
It is easy to see that for any two vectors $k,q\in\mathbb{R}^{d}$
and for any $k\in\mathbb{R}^{d}$

### v2

@@ A.2 Theoretical guarantees of SoftmaxBalance
A.3 MergeAndReduce
A.3.1 Pseudocode for MergeAndReduce
The pseudocode for MergeAndReduce is presented in 4.
@@ A.3.1 Pseudocode for MergeAndReduce
Figure 2: Illustration of the tree structure of MergeAndReduce
A.3.2 Theoretical guarantees of MergeAndReduce
We now present the full proof of Theorem 3.3.
@@ A.3.2 Theoretical guarantees of MergeAndReduce
Let us first consider the performance of the procedure at time steps which are multiples of $t$. Note that since in the statement of the theorem $T=\log_{2}(n/t)$, condition until in line while is redundant. Observe that at any such $j$-th step the procedure is an online implementation of the following simple offline recursive algorithm on dataset $\{\{k_{1},v_{1}\},\ldots,\{k_{j},v_{j}\}\}$:
@@
    Set $p=j/t$ and $i=1$. Split the dataset $\{\{k_{1},v_{1}\},\ldots,\{k_{j},v_{j}\}\}$ into batches $B^{0}_{1},\ldots,B^{0}_{p}$ of size $t$.
@@
    • Update $p\leftarrow\lfloor p/2\rfloor$, $i\leftarrow i+1$. Stop when $p=1$.
@@
Observe that after each iteration of step 2 the number of batches decreases by a factor of two. The maximum batch size is always bounded by $t$. This is because a batch $B_{l}$ at iteration $i$ of step 2 is a union of SoftmaxBalance$(B_{2l-1})$ and SoftmaxBalance$(B_{2l})$ for $B_{2l-1}$ and $B_{2l}$ at iteration $i-1$ of step 2, and SoftmaxBalance reduced the size of the dataset which it has been applied to at least by a factor of 2.
The memory of the procedure is the collection of memory cells $C^{i}$, and $|C^{i}|\leq t$. Since at time step $j$ at most $\log_{2}(p)\leq\log_{2}(n/t)$ memory cells are occupied, the total memory is bounded by $O(dt\log_{2}(n/t))=O(dtT),$ which, using the $\widetilde{O}$ notation, is equal to $\widetilde{O}(d\sqrt{d}e^{2r^{2}/\sqrt{d}}/{\varepsilon})$.
@@
$\leq\sum_{i=0}^{T-1}\left\|2^{i+1}\sum_{\{k,v\}\in B^{i+1}\cup C^{i+1}}\right.\left.\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-2^{i}\sum_{\{k,v\}\in B^{i}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v\right\|_{2}$
$\leq\sum_{i=0}^{T-1}2^{i}\left\|\sum_{\{k,v\}\in B^{i+1}\cup C^{i+1}}\right.\left.\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\in B^{i}\backslash(B^{i+1}\cup C^{i+1})}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v\right\|_{2}.$
We will refer to the $i$-th summand (starting from 0) on the right hand side as the error produced by the $i+1$-st iteration of step 2. At the $i+1$-st iteration of step 2 we apply SoftmaxBalance to $p/2^{i}$ batches $B^{i}_{1},B^{i}_{2},\ldots$ of size $t$, and we save the outputs of SoftmaxBalance in batches $C^{i+1},B^{i+1}_{1},B^{i+1}_{2},\ldots$. Therefore, by Theorem 3.2 and triangle inequality, the error vector produced by the procedure at the $i+1$-st iteration of step 2 has $l_{2}$ norm bounded by
$O\left(2^{i}\cdot\sqrt{s}\cdot\log(sn)\cdot\left(\frac{p}{2^{i}}\right)\cdot\ e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right)=O\left(\sqrt{s}\cdot\log(sn)\cdot\ p\cdot e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right),$
@@
$O\left(\sqrt{s}\cdot\log(sn)\cdot T\cdot p\cdot e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right).$
By definition, $p=j/t$. In order to ensure that the statement of the theorem is correct, the upper bound on the $l_{2}$ norm of the error vector of the procedure should be less than the desired error $e^{-r^{2}/\sqrt{d}}\cdot\max_{i\in[n]}\|v_{i}\|_{2}$:
$O\left(\sqrt{s}\cdot\log(sn)\cdot T\cdot\frac{j}{t}\cdot e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right)\leq\varepsilon j\cdot e^{-r^{2}/\sqrt{d}}\cdot\max_{j\in[n]}\|v_{j}\|_{2}.$
And, since by definition
@@
the above inequality holds.
Runtime during one time step. At worst, during $j$-th time step the algorithm has to launch SoftmaxBalance $\log_{2}(p)\leq\log_{2}(n/t)=T$ times on batches of size $t$, so the runtime is bounded by $O(dt^{2}T)$. In the $\widetilde{O}$ notation, the runtime is equal to $\widetilde{O}(d^{2}e^{4r^{2}/\sqrt{d}}/\varepsilon^{2}).$
@@
Appendix B Full Experimental Details
Experiments in Section 4.1 and Section 4.3 are performed on a single NVIDIA A100 GPU with 80GB VRAM, and the rest on a single NVIDIA RTX A6000 GPU with 48GB VRAM.
Implementation Detail.
B.1 Ablation Studies on Single Layer Attention Approximation
In this section we re-state with full details the single layer attention approximation experiments presented in Section 4.1.
We empirically evaluate the performance of BalanceKV for approximating a single attention layer, and compare it with independent uniform sampling. We use the pretrained $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] and TriviaQA dataset from LongBench [BLZ+23], and consider the $1^{st}$,$2^{nd}$ and $5^{th}$ layers of the models for attention approximation.
For given a prompt with length $n$, we store the corresponding query, key, and value embeddings for all layers. Denote a pair of embeddings in some layer by $(q_{1},k_{1},v_{1}),\ldots,(q_{n},k_{n},v_{n})$ and the goal is to approximate the attention $\text{Attn}(q_{j},K_{j},V_{j})$ for the latest $256$ queries, i.e. $j\in[n-256,n]$. Specifically, we keep several first and recent tokens separately and apply BalanceKV to the intermediate row vectors in $K_{j}$. This is motivated by StreamingLLM [XTC+23] as important contexts are likely contained in the first and latest tokens. We retain the first 256 embeddings and the recent ones from $n-256$ to $j$ and our compressed cache contains tokens whose indices are in $[256]\cup S\cup\{n-256,\dots,j\}$ where $S\subseteq[257,n-256]$ can be obtained from BalanceKV. We explore four compression parameters $T\in\{1,2,3,4\}$ which reduces the cache memory by a factor of $2^{-T}$. Let $z_{j}$ be our approximation using BalanceKV plus the recent and first few embeddings at the stream $j\in[n-256,n]$. We compute relative errors $\|z_{j}-\text{Attn}(q_{j},K_{j},V_{j})\|_{F}/\|\text{Attn}(q_{j},K_{j},V_{j})\|_{F}$ for all $j\in[n-256,n]$, batches, heads and input prompts in the dataset. We repeat this with 10 different random seeds and compute their average and standard deviations. We also compare our method to independent uniform sampling, in which we replace the application of BalanceKV with sampling a $2^{-T}$ fraction of key and value embeddings with indices in $[257,n-256]$ uniformly at random. The results are reported in Figure 1.
Next we present the results of the ablation studies described in Section 4.1 which demonstrate how batch size and compression rate affect the relative error in attention approximation for layers 1 and 15 for $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$.
Batch Size 1/2 1/4 1/8
256 0.0603 0.1190 0.1793
128 0.0320 0.0624 0.0922
64 0.0189 0.0349 0.0508
(a) Layer 1 Runtime (s)
1/2 1/4 1/8
256 0.1036 0.1764 0.2655
128 0.1082 0.1833 0.2741
64 0.1137 0.1921 0.2858
(b) Layer 1 Relative Error
1/2 1/4 1/8
256 0.3920 0.4505 0.5096
128 0.3654 0.3951 0.4256
64 0.3592 0.3753 0.3910
(c) Layer 15 Runtime (s)
1/2 1/4 1/8
256 0.1107 0.1935 0.2798
128 0.1121 0.1952 0.2813
64 0.1141 0.1978 0.2845
(d) Layer 15 Relative Error
Figure 3: Runtime and relative error for across different layers and block sizes. In each figure the rows are corresponding to various batch sizes and columns corresponding to various compression rates
B.2 End-to-End Evaluation on LongBench
We now provide the complete experimental details on the end-to-end evaluation in Section 4.2. We benchmark our algorithm on LongBench dataset [BLZ+23], a comprehensive collection of datasets designed to evaluate the long-context understanding capabilities of large language models. Specifically, we test a version of uniform length distribution (LongBench-E). The benchmark consists of various long-text application scenarios, including single-document question-answering, multi-document question-answering, summarization, few-shot learning, synthetic tasks and code completion. We use BalanceKV to compress the key value cache generated in the prefill stage, and maintain all streamed embeddings $(q_{j},k_{j},v_{j})$ during the token decoding/generation stage. This is because the number of generated tokens is much smaller than the input sequence length. We set $b=256$ and $T=2$, achieving a consistent compression rate of $0.25$ across all inputs.
We evaluate our method against several token-level key value cache compression schemes, including
StreamingLLM  [XTC+23], SnapKV [LHY+24], and PyramidKV [CZG+24] as well as uniform sampling described in Section 4.1. We use their implementations from MInference [JLZ+24], and configure their hyperparameters to match a uniform compression rate with $0.25$. We follow the same evaluation metrics from [BLZ+23]. We test them on $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ as well as a bigger 14B and 32B parameter models $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{14B}$-$\mathtt{Instruct}$ and $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{32B}$-$\mathtt{Instruct}$ [YYH+24, Tea24d], with results summarized in Table 1.
Our method consistently achieves the highest average performance among compression methods and across all models, demonstrating its effectiveness in preserving model quality for the cache compression. Notably, on the triviaqa dataset, it achieves near-exact scores compared to uncompressed baselines (e.g., 80.68 vs. 81.14 with Qwen2.5-32B), highlighting its ability to retain high-quality information. We observe that uniform sampling performs competitively with our method and this result justifies that a subset obtained from discrepancy theory has practical impacts on various LLM tasks.
B.3 Needle-In-A-Haystack
In this section we report the plots corresponding to the Needle in a Haystack experiment described in Section 4.3. They are presented in Figure 4.
Figure 4: Comparison of performance on Needle in a Haystack task using $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$. The methods corresponding to figures from top to bottom are StreamingLLM, SnapKV, PyramidKV, Unif. Sampling and BalanceKV respectively.
B.4 System Efficiency Metrics
In this section we present the prefill and decoding time numbers in Table 2 as described in the system efficiency experimental details in Section 4.4.
Method Prefill Time (sec) Decoding Time (sec)
Exact 3.032 37.769
SnapKV 3.755 40.426
PyramidKV 3.748 37.241
StreamingLLM 3.681 40.276
BalanceKV 3.662 38.054
Table 2: Minimum wall-clock runtime (in seconds) over 10 trials for prefill and decoding stages.
Appendix C Lower bound
[…]
The proof will be a reduction to the well-known INDEX problem.
C.1 Reduction to the INDEX problem
Definition C.1 (The INDEX problem).
Alice gets a bit string $x\sim\text{Unif}\{0,1\}^{n}$ and Bob gets $i\sim\text{Unif}[n]$. Then, the goal is to compute $f(x,i)=x_{i}$ on Bob’s end with a single message $m$ from Alice. Denote by $R_{\delta}^{pub,\to}$ the public coin one-way communication complexity of computing a function $f(x,y)$ with error probability at most $\delta$: Alice holds $x$, Bob holds $y$, they share a source of random bits and Alice sends a single message to Bob, after which he must output the correct answer with probability at least $1-\delta$.
Theorem C.2 (Proven in [Woo20]).
$R_{2/3}^{pub,\to}(INDEX)\geq\Omega(n).$
@@ C.1 Reduction to the INDEX problem
Let $c$ be the small constant such that $R_{2/3}^{pub,\to}(INDEX)\geq c\cdot n$.
@@
The reduction.
Suppose Alice’s input to the INDEX problem is a bit string $x\in\{0,1\}^{n}$ of length $n=\min\{\frac{1}{\varepsilon^{2}},d\exp(2r^{2}/\sqrt{d})\}$. Using public coins, Alice and Bob jointly generate $n/d$ key vectors $\tilde{k}_{1},\ldots,\tilde{k}_{n/d}\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d},$ function $\pi:[n]\to[n/d]\times[d]$ which randomly partitions the $n$ bits into groups of size $d$, and $n$ random signs $\sigma_{1},\ldots,\sigma_{n}\sim\text{Unif}\{-1,1\}$.
Let $\pi(i)_{1}\in[n/d]$ be the first component of $\pi(i)$ and $\pi(i)_{2}\in[d]$ – the second component of $\pi(i)$. Let $e_{1},\ldots,e_{d}$ be the standard orthonormal basis in $\mathbb{R}^{d}$. We build the dataset of $n$ key-value pairs in the following way. We associate with the $i$-th bit the key vector $k_{i}\coloneqq\tilde{k}_{\pi(i)_{1}}$ and the value vector $v_{i}\coloneqq\sigma_{i}\cdot e_{\pi(i)_{2}}$.
Define
$U=\{\{k_{i},v_{i}\}:x_{i}=1\}$
the set of key-value pairs corresponding to entries 1 in Alice’s bit string $x$. Alice instantiates the streaming algorithm for approximating $\text{Attn}(\cdot,K,V)$ with space complexity $\frac{c}{C}\cdot\min\{\frac{1}{\varepsilon^{2}},d\exp(2r^{2}/\sqrt{d})\}$ with a big enough constant $C$ which we specify later and sends the state of it’s memory before reading $q$ to Bob. When Bob receives the message, he uses it to approximate $\text{Attn}(q_{i},K,V)$ where $q_{i}=k_{i}$. If the value written in the only non-zero coordinate of $v_{i}$ is larger than $\frac{1}{40}\cdot\frac{\exp(r^{2}/\sqrt{d})}{\max\{d\exp(r^{2}/\sqrt{d}),|U|\}}$, Bob reports that the $i$-th bit of Alice’s string is equal to 1, and otherwise  - 0.
Analysis of the reduction.
Proof sketch.
Before moving to formal proofs we briefly outline the main idea of the analysis. Observe that, by the choice of key vectors, any $\exp(\langle k_{j},q_{i}\rangle/\sqrt{d})$ for $j\neq i$ is in expectation insignificantly small compared to $\exp(\langle k_{i},q_{i}\rangle/\sqrt{d})$  – sometimes we will even refer to these terms as “noise”. This statement is formalized in Lemma C.4. Therefore, when Bob computes an approximation to $\text{Attn}(q_{i},K,V)$, he will observe a large value in the coordinate where $v_{i}$ is non-zero if $\{k_{i},v_{i}\}\in U$ and a small value otherwise.
Lemma C.3.
@@
$\mathbb{E}_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}[\exp(C\langle x,y\rangle/\sqrt{d})]=\Theta(1)$ for any constant $C$.
Proof.